Percent Composition and Empirical Formula





1. Phosphoric acid (H3PO4) is a colorless, syrupy liquid used in detergents, fertilizers, toothpastes,




and in carbonated beverages for a “tangy” flavor. Calculate the percent composition by mass of H, P,




and O in this compound.





2. Propane is hydrocarbon, a compound composed only of carbon and hydrogen. It is 81.82%




carbon and 18.18% hydrogen. What is the empirical formula?

1
Expert's answer
2022-03-13T15:20:38-0400

1. Molecular mass of H3PO4 = 3×1+31+64=98

% mass of H = "\frac{3}{98}= 3.06 %" %


% mass of P = "\frac{31}{98}= 31.63 %" %


% mass of O = "\frac{64}{98}= 65.306" %


2.

18.18 % of H present in 81.82 % of C


18.18 g of H present in 81.82 g of C


Mole of H present "\frac{18.18}{18.18}= 18"mole

Mole of C =6.81

Emprical formula = CH3